OpenStudy (anonymous):

solve the system of equations an choose the correct ordered pair 6x-8y+6=36 6y=4x-24

OpenStudy (anonymous):

6x-8y+6=36\\ 6y=4x-2\\ 6 x -8 y = 30\\ -2 x +3y =-1\\ 6 x -8 y = 30\\ -6 x +9y =-3\\

OpenStudy (anonymous):

Add them y= 27 Replace y by 27 in one of the equations and you get x=41

OpenStudy (anonymous):

(41, 27) is the ordered pair solution.

OpenStudy (anonymous):

Thanks.!